A wire resistance 18 ohm is drawn until it's radius reduce 1\2 of its original radius then resistance of the wire is
1
Expert's answer
2019-07-19T14:12:50-0400

"R_{1}" = 18 ohm

Relation to resistivity and conductivity: "R_{1}=\rho\frac{l}{S_{1}}"

"r_{2}=\frac{1}{2}r_{1}" , "S_{1}=\pi {r_{1}^2}"

and "S_{2}=\pi {r_{2}^2} = \frac{1}{4}{S_1}=\frac{1}{4}{\frac{\rho{l}}{R_{1}}}"

Then "R_{2}=\rho\frac{l}{S_{2}} = \frac{4{R_{1}{\rho}l}}{{\rho}l}=4R_{1}=72 ohm"